# Salem numbers less than the plastic constant

### v1

@@ Abstract
A list of Salem numbers less than $1.3$ is available on M. Mossinghoff’s website ([2]). This list is certified complete up to degree $44$ in [9], and it includes only one Salem number of degree $46$. The objective of the present work is to advance the understanding of Salem numbers by extending the list [2] through the provision of a list of Salem numbers less than the plastic constant, denoted by $\eta$, which is approximately equal to $1.324718$. The algorithmic approach used is based on Integer Linear Programming.
@@
Keywords— Salem numbers, plastic constant, reciprocal polynomials
@@ 1 Introduction
In the following lines, we recall some basic elements related to Salem numbers.
@@ A Salem number is an algebraic integer $\tau$ such that:
  • All of its Galois conjugates (i.e., the other roots of its minimal polynomial) have an absolute value less than or equal to $1$,
  • At least one of its Galois conjugates has an absolute value equal to $1$.
@@
The Salem number $\tau_{0}=1.176280\dots$ has the minimal polynomial $P_{0}(x)=x^{10}+x^{9}-x^{7}-x^{6}-x^{5}-x^{4}-x^{3}+x+1$, known as the Lehmer polynomial. The real number $\tau_{0}$ is the smallest known Salem number to date (D. Lehmer, 1933).
@@ 1.2 Some properties
  • The minimal polynomial of a Salem number $\tau$ is reciprocal and has an even degree greater than or equal to 4.
@@
  • All the roots of the minimal polynomial of $\tau$, except for $\tau$ and $\frac{1}{\tau}$, have an absolute value equal to $1$.
@@ 1.3 Small Salem numbers
Salem numbers less than $1.3$ are referred to as small in the literature. In 1977, D. Boyd provided a list of $39$ such numbers in [5]. The following year, D. Boyd added $4$ more small Salem numbers in [6], before $4$ additional small Salem numbers were discovered and published by M. Mossinghoff in 1993 ([7]). Notably, among the $4$ values provided by M. Mossinghoff, one has a degree of $46$, which is the highest known degree for a small Salem number.
The compilation of these sources thus yields a list of $47$ small Salem numbers, which is available on M. Mossinghoff’s website ([2]). This list provides, for each Salem number, its degree and the coefficients of its minimal polynomial. Naturally, only the relevant part of the coefficients is given, as these polynomials are reciprocal.
@@
In this work, we propose to extend the list [2] by providing Salem numbers that are smaller than the plastic constant $1.324718\ldots$. The choice of the plastic constant as the threshold may seem somewhat arbitrary, but in fact the aim is to go slightly beyond $1.3$ in order to advance our understanding of Salem numbers by compiling an extended list. As a reminder, the plastic constant is the unique real root of the polynomial $x^{3}-x-1$. This constant is also the smallest Pisot number, where a Pisot number is a real algebraic integer strictly greater than $1$ whose Galois conjugates all have absolute values strictly less than $1$. The plastic constant plays an important role in number theory. For example, C. Smyth shows in [1] that if the Mahler measure of a polynomial in $\mathbb{Z}[x]$ is less than this constant, then the polynomial must be reciprocal.
To compile our extended list, we will use a variation of a method based on integer linear optimization, which has allowed us to obtain interesting results in various contexts. Notably, we followed this approach to provide, in [10], a list of degree $16$ irreducible monic polynomials with integer coefficients, all real and strictly positive roots and minimal trace (specifically, with trace $29$). Similarly, we provided examples of Salem numbers of degree $34$ and trace $-3$ in [11] and [12].
Here, it is not the trace of the Salem number that interests us, but its size. Our approach, based on integer linear optimization, is well-suited for addressing this problem. In the following lines, we only outline the method, as it is described in detail in the aforementioned articles.
@@ 2.1 Reduction to a half-degree problem
where the $C_{i}$ are the Chebyshev polynomials associated with the interval $(-2,2)$. As a reminder, we have
@@ 2.2 Reformulation as an Optimization Problem
In practice, once a $(d-2)$-tuple $(\beta_{1},\dots,\beta_{d-2})$ is drawn, we solve the problem. If no $d$-tuple $(a_{0},\dots,a_{d-1})$ is found that satisfies all the constraints, we immediately draw another sample. Conversely, if such a $d$-tuple $(a_{0},\dots,a_{d-1})$ is found, this means we have found a polynomial $q$ (of degree $d$) of the type $Q$, which allows us to reconstruct a Salem polynomial $P$ of the desired type. If $q$ is irreducible in $\mathbb{Z}[x]$, we add the polynomial $P$ reconstructed from $q$ to our list. If $q$ is not irreducible, we keep the factor of $q$, say $q_{1}$, that has a root between $2$ and $\eta+\displaystyle\frac{1}{\eta}$, and we reconstruct a Salem polynomial from $q_{1}$. The following flowchart summarizes the approach used:
@@
3 Implementation and results
3.1 Implementation
@@ 3.2 Results
The implementation of the algorithmic approach presented in this work did not lead to the discovery of new Salem numbers smaller than $1.3$. In the following table, we therefore only present the Salem numbers between $1.3$ and $\eta$ resulting from our calculations. To make the table easy to read, we follow the presentation chosen in [2], in which the different columns respectively show the degrees of the Salem numbers, the numbers themselves, and the relevant coefficients of the minimal polynomials (since these are reciprocal)
@@
20 1.324351296267 1 -1 0 -1 0 1 0 0 0 -1 1
24 1.324441583513 1 0 0 0 -1 -1 -1 -1 -1 0 0 0 1
30 1.324509572546 1 0 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1
32 1.324560796779 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
30 1.324599405908 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1 -1 0 0
30 1.324628516131 1 0 -1 -1 0 1 0 -1 -1 0 1 0 -1 -1 0 1
34 1.324650470079 1 0 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1
34 1.324667030364 1 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1
32 1.324679524104 1 0 0 -1 -1 -1 0 0 1 0 0 -1 -1 -1 0 0 1
30 1.324688951068 1 0 0 0 0 0 -1 -1 -1 -1 -2 -2 -2 -2 -2 -3
32 1.324696064745 1 -2 1 0 0 0 0 -1 1 0 0 -1 1 -1 1 0 -1
40 1.324701433212 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
40 1.324705484871 1 0 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0
36 1.324708542862 1 0 0 0 -1 -1 -1 -1 -1 0 0 0 1 0 0 0 -1 -1 -1
36 1.324710850968 1 0 0 0 0 -1 -1 -1 -1 -2 -1 -1 -1 -1 0 0 0 0 1
40 1.324712593127 1 0 -1 -1 0 1 0 -1 -1 0 1 0 -1 -1 0 1 0 -1 -1 0 1
38 1.324713908140 1 -1 0 0 -1 0 1 -1 1 0 -1 0 0 -1 1 0 0 0 0 -1
46 1.324714900752 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
46 1.324715650017 1 0 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1
46 1.324716215600 1 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1
44 1.324716642533 1 -1 0 -1 1 0 0 -1 0 0 1 -1 0 -1 1 0 0 -1 0 0 1 -1 0
50 1.324716964809 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
48 1.324717208083 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1
52 1.324717391722 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
48 1.324717530347 1 0 0 -1 -1 -1 0 0 1 0 0 -1 -1 -1 0 0 1 0 0 -1 -1 -1 0 0 1
44 1.324717634990 1 -1 0 0 -1 1 -1 0 0 0 1 -1 1 -1 0 1 -2 1 -1 0 1 -1 1
54 1.324717713983 1 0 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1 0 -1
56 1.324717773612 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
54 1.324717818625 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1 -1 0 0 0 -1 1 -1 0 0
58 1.324717852604 1 1 0 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1 -1
54 1.324717878254 1 -1 0 -1 1 0 0 -1 0 0 1 -1 0 -1 1 0 0 -1 0 0 1 -1 0 -1 1 0 0 -1
58 1.324717897616 1 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1 0 0 -1
50 1.324717912232 1 0 0 0 -1 -1 -1 -1 0 -1 0 0 -1 0 0 0 1 0 1 0 -1 0 -1 -1 0 -1
60 1.324717946211 1 0 0 0 -1 -1 -1 -1 -1 0 0 0 1 0 0 0 -1 -1 -1 -1 -1 0 0 0 1 0 0 0 -1 -1 -1
@@ 4 Conclusion
A new Salem number smaller than $1.3$ would of course have been welcome. In any case, our line of research has proven fruitful, as dozens of Salem numbers lower than the plastic constant are now available. It should be noted, of course, that this algorithmic approach does not guarantee the exhaustiveness of the tables it provides. But since it is based on random sampling, the repeated discovery of values already present in our list suggests that any missing values are likely to be few. In particular, the repeated discovery of known Salem numbers below 1.3 using this algorithmic approach, without finding new ones, suggests that List [2] might be complete well beyond degree $44$ certified in [9].

### v3

# Salem numbers less than $49/37$

@@ Abstract
A certain number of lists of small Salem numbers, some of which are certified as complete, are available online. Notably, the website of M. J. Mossinghoff features a list of 47 Salem numbers smaller than 1.3 ([3]), as well as complete lists of Salem numbers of fixed degrees that are smaller than various bounds ([2]). The objective of this work is to advance the understanding of Salem numbers by providing a list of Salem numbers smaller than a threshold of $49/37$ ($\approx 1.324324$) through the implementation of a method based on integer linear programming and uniform sampling of root separators. Beyond the intrinsic interest of the newly detected Salem numbers, the rediscovery of already known Salem numbers through this alternative method could offer valuable insights into the potential completeness of already existing lists for degrees where it has not yet been proven.
@@
Keywords— Salem numbers, Mahler measure
@@ 1 Introduction
Salem numbers have drawn significant interest since at least the 1940s. The focus on small Salem numbers has been particularly prominent, partly due to their connection with Lehmer’s problem, which concerns integer polynomials with small Mahler measure. Numerous researchers have investigated these numbers, with efforts tracing back at least to D. W. Boyd’s work in the 1970s. In the following lines, we provide a brief overview of some key properties of Salem numbers.
@@ A Salem number is an algebraic integer $\tau$ such that:
  • All of its Galois conjugates (i.e., the other roots of its minimal polynomial) have absolute value less than or equal to $1$,
  • At least one of its Galois conjugates has absolute value equal to $1$.
@@
The Salem number $\tau_{0}=1.176280\dots$ has the minimal polynomial $P_{0}(x)=x^{10}+x^{9}-x^{7}-x^{6}-x^{5}-x^{4}-x^{3}+x+1$, known as the Lehmer polynomial. The real number $\tau_{0}$ is the smallest known Salem number to date ([18]).
@@ 1.2 Some properties
  • The minimal polynomial of a Salem number $\tau$ is reciprocal and has even degree greater than or equal to 4.
@@
  • All the roots of the minimal polynomial of $\tau$, except for $\tau$ and $\frac{1}{\tau}$, have absolute value equal to $1$.
@@ 1.3 Small Salem numbers
Salem numbers less than $1.3$ are referred to as small in the literature. In 1977, D. W. Boyd provided a list of $39$ such numbers in [16]. The following year, D. W. Boyd added $4$ more small Salem numbers in [7], before $4$ additional small Salem numbers were discovered and published by M. J. Mossinghoff in 1993 ([8]). Notably, among the $4$ values provided by M. J. Mossinghoff, one has a degree of $46$, which is the highest known degree for a small Salem number.
The compilation of these sources thus yields a list of $47$ small Salem numbers, which is available on M. J. Mossinghoff’s website ([3]). This list provides, for each Salem number, its degree and the coefficients of its minimal polynomial. Naturally, only the relevant part of the coefficients is given, as these polynomials are reciprocal.
@@
In this work, we propose to extend the list [3] by providing Salem numbers that are smaller than $\eta=49/37$. The value chosen for $\eta$ might seem somewhat arbitrary and therefore deserves a brief justification. Let us point out that the real number $\eta$ is close to the plastic constant, the real root of $x^{3}-x-1$, whose approximate value is $1.324718$. The plastic constant holds particular significance in the study of Salem numbers. Let us recall that a Pisot-Vijayaraghavan number is a real algebraic integer greater than $1$, all of whose Galois conjugates are less than $1$ in absolute value. R. Salem provided a construction which, given a Pisot-Vijayaraghavan number, produces two sequences of Salem numbers, one approaching the Pisot-Vijayaraghavan number from below and the other from above ([15], [5]). In turn, D. W. Boyd proved that every Salem number appears as a member of one of the sequences of Salem numbers arising from Pisot-Vijayaraghavan numbers (see [16] or [5]). It is worth mentioning that the plastic constant, which is the smallest Pisot-Vijayaraghavan number, is the smallest known limit point of the Salem numbers. From the previous lines, we deduce that there is an infinite number of Salem numbers smaller than the plastic constant, and it is possible to explicitly construct a sequence of such numbers. This explains why the value $1.3$ has traditionally been chosen as the threshold for defining small Salem numbers. The choice in this work is to set a new threshold exceeding $1.3$, while remaining strictly below the plastic constant, for clear reasons mentioned above. Our choice fell on the value $49/37$, which has an approximate value of $1.324324$.
To construct our extended list, we used a variation of a method based on integer linear optimization, which has allowed us to obtain interesting results in various contexts. Notably, we followed this approach to provide, in [11], a list of degree $16$ irreducible monic polynomials with integer coefficients, all real and strictly positive roots and minimal trace (specifically, with trace $29$). Similarly, we provided examples of Salem numbers of degree $34$ and trace $-3$ in [12] and [13].
Here, it is not the trace of Salem numbers that interests us, but their size. Our approach, based on integer linear optimization, is well-suited for addressing this problem. In the following lines, we only outline the method, as it is described in detail in the aforementioned articles.
@@ 2.1 Reduction to a half-degree problem
where the $C_{i}$’s are the Chebyshev polynomials associated with the interval $(-2,2)$. As a reminder, we have
@@ 2.2 Reformulation as an Optimization Problem
2.3 Practical implementation
In practice, the $\beta_{i}$ points are drawn uniformly from $(-2,2)$ in each trial. Once a $(d-2)$-tuple $(\beta_{1},\dots,\beta_{d-2})$ is drawn, we solve the problem. If no $d$-tuple $(a_{0},\dots,a_{d-1})$ is found that satisfies all the constraints, we immediately draw another sample. Conversely, if such a $d$-tuple $(a_{0},\dots,a_{d-1})$ is found, this means we have found a polynomial $q$ (of degree $d$) of the type $Q$, which allows us to reconstruct a Salem polynomial $P$ of the desired type. If $q$ is irreducible in $\mathbb{Z}[x]$, we add the polynomial $P$ reconstructed from $q$ to our list. If $q$ is not irreducible, we keep the factor of $q$, say $q_{1}$, that has a root between $2$ and $\eta+\displaystyle\frac{1}{\eta}$, and we reconstruct a Salem polynomial from $q_{1}$. The following flowchart summarizes the approach used:
@@ 2.3 Practical implementation
We performed calculations related to the search for degree $d$ polynomials with real roots for all values of $d$ between $10$ and $32$. From the perspective of Salem polynomials, this means we searched for such polynomials up to degree $642d=$ 64.
For small values of $d$, we carried out $800,000$ uniform random draws (this number was chosen because we had access to an $80$-core server), and we progressively increased the number of draws to $1,600,000$ for larger values of $d$.
3 Technical details and results
3.1 Technical details
@@ 3.2 Results
The implementation of the algorithmic approach presented in this work found all known Salem numbers less than $1.3$, but no new numbers less than this value. In the following table, we therefore only present the Salem numbers between $1.3$ and $\eta$ resulting from our calculations. To make the table easy to read, we follow the presentation chosen in [3], in which the different columns respectively show the degrees of the Salem numbers, the numbers themselves, and the relevant coefficients of the minimal polynomials (since these are reciprocal). Up to degree $24$, the Salem numbers found using our method are already known and are listed on M. J. Mossinghoff’s website ([3]). These polynomials appeared in a $2014$ article by K. G. Hare and M. J. Mossinghoff ([17] ), in a study of specific representations of the sequences of Salem numbers associated with small Pisot-Vijayaraghavan numbers. As for us, we provide 10 new Salem numbers less than $49/37$, with degrees ranging from $26$ to $44$. They are highlighted in bold in the table above.
@@
The fact that all known Salem numbers less than 1.3, as well as all Salem numbers less than $49/37$ with a degree smaller than 24, have been recovered using our method seems highly significant to us. This is not only a testament to the effectiveness of our approach but also provides a basis for assessing the completeness of the lists available online. Indeed, two factors must be considered. The first is that our method is based on uniform random sampling, meaning it is free from any bias in its ability to detect values. The second is that not only were the already known values rediscovered using this method, but they were also found repeatedly, many times over. Similarly, the newly discovered values appeared numerous times throughout the program’s execution. Thus, while this is not theoretically proven, there are reasonable grounds to believe that for the degrees examined, we have likely identified all that could be found. The same argument leads us to think that the list of Salem numbers less than 1.3, certified complete up to degree 44, is likely complete up to significantly higher degrees, given that all its values were repeatedly detected without any new ones emerging.
@@ 4 Conclusion
A new Salem number smaller than $1.3$ would of course have been welcome. In any case, our line of research has proven fruitful, as $10$ new Salem numbers lower than $49/37$ are now available. It should be noted, of course, that this algorithmic approach does not guarantee the exhaustiveness of the table it provides. But since it is based on random sampling, the repeated discovery of values already present in our list suggests that any missing values are likely to be few. In particular, the repeated discovery of known Salem numbers below $1.3$ using this algorithmic approach, without finding new ones, suggests that List [3] might be complete well beyond degree $44$ certified in [10].
5 Acknowledgement
The author of this article wishes to express deep gratitude to the anonymous referee, whose valuable advice has greatly improved the final manuscript. They are sincerely thanked for it.